# Distributed Primal-Dual Algorithms: Unification, Connections, and Insights

### v1

@@ Abstract
We study primal-dual algorithms for general empirical risk minimization problems in distributed settings, focusing on two prominent classes of algorithms. The first class is the communication-efficient distributed dual coordinate ascent (CoCoA), derived from the coordinate ascent method for solving the dual problem. The second class is the alternating direction method of multipliers (ADMM), including consensus ADMM, linearized ADMM, and proximal ADMM. We demonstrate that both classes of algorithms can be transformed into a unified update form that involves only primal and dual variables. This discovery reveals key connections between the two classes of algorithms: CoCoA can be interpreted as a special case of proximal ADMM for solving the dual problem, while consensus ADMM is closely related to a proximal ADMM algorithm. This discovery provides the insight that by adjusting the augmented Lagrangian parameter, we can easily enable the ADMM variants to outperform the CoCoA variants. We further explore linearized versions of ADMM and analyze the effects of tuning parameters on these ADMM variants in the distributed setting. Our theoretical findings are supported by extensive simulation studies and real-world data analysis.
Machine Learning, ICML
@@ 1 Introduction
In this paper, we consider algorithms solving a federated learning problem where $K$ machines collaboratively solve a general empirical risk minimization problem, based on $n$ samples $\left\{x_{i}\right\}_{i=1}^{n}\subset\mathbb{R}^{d}$ distributed in all machines,
$\underset{w\in\mathbb{R}^{d}}{\min}\left\{\mathcal{P}(w):=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{P}_{k}}\ell_{i}(w^{\top}x_{i})+g(w)\right\}.$ (1)
In this formulation, $\left\{\mathcal{P}_{k}\right\}_{k=1}^{K}$ denotes the partition of these samples in machines, $|\mathcal{P}_{k}|=n_{k}$ and $\sum_{k=1}^{K}n_{k}=n$. The parameter of common interest is $w\in\mathbb{R}^{d}$, $\ell_{i}(\cdot)$ is the local convex loss function potentially incorporating label information for the $i$-th data sample, and $g(\cdot)$ is a convex regularization penalty possibly non-smooth. Examples of loss functions include least squares loss, quantile loss, Huber loss, and hinge loss, while examples of penalty functions include the $\ell_{1}$ norm, $\ell_{2}$ norm and elastic net. Problems of this type is pervasive in various applications of machine learning.
@@
Over the past decades, efficient distributed algorithms have been designed to solve problem (1). One popular class of the algorithms to solve the problem (1) is communication-efficient distributed dual coordinate ascent (CoCoA) framework and its variants (Yang, 2013; Ma et al., 2015; Smith et al., 2018; Jaggi et al., 2014; Dünner et al., 2018; Lee & Chang, 2020). This approach originates from adapting the dual coordinate descent method (Shalev-Shwartz & Zhang, 2013) for distributed scenarios. It create an upper bound of the dual objective which enables parallel update of the corresponding dual variables on individual machines, under the assumption that the penalty function of $g$ is strongly convex.
@@
Another class of methods that solves (1) is the alternating direction method of multipliers (ADMM) and its variants (Boyd et al., 2011). For example, consensus ADMM method has been recognized as an effective way to solve federated learning problems, which tailors the classical two-block ADMM method for distributed scenarios through duplicating many local variables. Recently, significant research has focused on developing ADMM variants to address the federated learning problems. An extension is generalized (or proximal) ADMM algorithms, for example (Deng & Yin, 2016), to achieve faster convergence rate for the ADMM algorithm. The reviews (Yang et al., 2022; Han, 2022; Glowinski, 2014; Maneesha & Swarup, 2021) provide comprehensive discussions on the recent advancements of using ADMM variants in distributed optimization.
@@
In this study, we present an interesting discovery that CoCoA, consensus ADMM, and two distributed proximal ADMM algorithms (Deng & Yin, 2016; Deng et al., 2017) we identified can all be cast into the same kind of update rules that only involve the primal variables and dual variables. Up to our knowledge, this is an original discovery, and the unified update rules result in (1) an easy calculation of the dual gap, (2) a novel understandings on the connections among CoCoA and ADMM-type algorithms in the existing literature, and (3) a unified proof for the convergence of ADMM-type algorithms. A major outcome is that the update rule of CoCoA is identical to a special proximal ADMM algorithm with a specific selection of step size, whereas this proximal ADMM algorithm has a strong connection with a consensus ADMM algorithm.
@@
    We reformulate consensus ADMM, linear consensus ADMM, two proximal ADMM algorithms, and CoCoA into a unified update form that involves only the primal variable $w$ and the dual variable $v$. The two proximal ADMM algorithms are original to solves the regularized federated learning problem (1). We also propose a primal-dual gap difference certificate to serve as a stopping criterion for consensus ADMM.
@@
    Based on this unified formulation, we establish connections among these three distributed algorithm frameworks. The dual variable updates in the CoCoA framework is equivalent with the first proximal ADMM algorithm, which in turn associated with the consensus ADMM. The linearized version of the consensus ADMM is associated with the second proximal ADMM formulation that enable leveraging the closed-form expression of the proximal operators of the loss.
@@
    We thoroughly study the effects of the tuning parameters in both the proximal ADMM and consensus ADMM and use synthetic and real data experiments to verify our results.
The remaining part of the article is organized as follows. In Section 2, we introduce the five algorithms and cast them into a unified form of primal-dual update. In Section 3, we use the unified update form to evaluate the connections between algorithms. In Section 4, further provides a unified proof of the convergence for all algorithms, leveraging the update form. Section 5 use experiments to validates our results, and Section 6 concludes the article. All proofs are included in the appendix.
@@
Let $X=[x_{1},\ldots,x_{n}]\in\mathbb{R}^{d\times n}$ represent the whole training data matrix. We use a column vector $v=[v_{1},\ldots,v_{n}]^{\top}\in\mathbb{R}^{n}$ to denote the dual variable. We use $v_{[k]}\in\mathbb{R}^{n_{k}}$ and $X_{[k]}\in\mathbb{R}^{d\times n_{k}}$ to denote the dual variable block and training data block on the $k$-th machine, where $X=[X_{[1]},\ldots,X_{[K]}]$ and $v=(v_{[1]}^{\top},\ldots,v_{[K]})^{\top}$. Correspondingly, we define $\ell_{[k]}^{*}(v_{[k]})=\sum_{i\in\mathcal{P}_{k}}\ell_{i}^{*}(v_{i})$. For a positive semidefinite matrix $S$, we denote by $\|x\|_{S}:=\sqrt{x^{\top}Sx}.$ $\lambda_{\max}(M)$ denotes the largest eigenvalue of matrix $M$. The proximal operator for any convex function $f(\cdot)$ and positive constant $\lambda>0$ is defined as
$\text{prox}_{\lambda f}(v)=\arg\min_{x\in\mathbb{R}^{m}}\left(f(x)+\frac{1}{2\lambda}\|x-v\|^{2}\right).$
@@ 2 Distributed Algorithms via Primal and Dual Updates
In this section, we demonstrate that various distributed algorithms, including CoCoA algorithm with ridge penalty (Jaggi et al., 2014; Ma et al., 2015, 2021; Smith et al., 2018), global consensus ADMM algorithm with regularization (Boyd et al., 2011) and its linearized variant (Lin et al., 2011), and two proximal ADMM (Deng & Yin, 2016) that enable distributed solution of problem (1), can all be cast into an update form that only involves the updates for the primal and the dual variables. As we will see in the next section, the transformations will reveal valuable connections between the different techniques. The transformation of the CoCoA algorithm and the proximal ADMM algorithms will both rely on the dual problem of problem (1) below (see Appendix A).
$\underset{v\in\mathbb{R}^{n}}{\min}\left\{\mathcal{D}(v):=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(v_{i})+g^{*}\left(-Xv/n\right)\right\}.$ (2)
Here $f^{*}$ denotes the Fenchel conjugate of a proper, convex, and lower semi-continuous function $f:\mathbb{R}^{m}\to\mathbb{R}$, defined as $f^{*}(u)=\sup_{v\in\mathbb{R}^{m}}\left\{\langle u,v\rangle-f(v)\right\}$.
2.1 CoCoA with Ridge Penalty
With the ridge penalty $g(w)=\lambda\left\lVert w\right\rVert_{2}^{2}/2$, the dual form (2) can be written as
$\underset{v\in\mathbb{R}^{n}}{\min}\left\{\mathcal{D}(v):=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(v_{i})+\frac{1}{2n^{2}\lambda}\|Xv\|^{2}\right\}.$
CoCoA follows a majorization-minimization scheme and minimizes $Q(v|v^{(t)})$, an upper bound function of $\mathcal{D}(v)$ in each iteration $t$ with the current dual variable $v^{(t)}$ (Smith et al., 2018) in a distributed manner:
$\begin{split}&Q(v|v^{(t)})=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(v_{i})+\frac{1}{2n^{2}\lambda}\Big{\|}\sum_{k=1}^{K}X_{[k]}v_{[k]}^{(t)}\Big{\|}^{2}\\ &+\frac{1}{n^{2}\lambda}\langle X^{\top}Xv^{(t)},v\rangle+\frac{\sigma}{2n^{2}\lambda}\sum_{k=1}^{K}\Big{\|}v_{[k]}-v_{[k]}^{(t)}\Big{\|}_{X_{[k]}^{\top}X_{[k]}}^{2}.\end{split}$
The specific algorithm can be expressed as
$\begin{split}\tilde{v}^{(t)}&=\underset{v\in\mathbb{R}^{n}}{\arg\min}\;Q(v|v^{(t)}),\\ v^{(t+1)}&=v^{(t)}+\gamma(\tilde{v}^{(t)}-v^{(t)}),\\ w^{(t+1)}&=-\frac{1}{n\lambda}\sum_{k=1}^{K}X_{[k]}v_{[k]}^{(t+1)}.\end{split}$
where the $w$-step recovers the primal decision variable, according to the KKT conditions, while the $v$-step aims to decrease $\mathcal{D}(v)$. The tuning parameter $\sigma$ governs the approximation of the original dual problem and $\gamma$ controlling the adjustment of the dual variable. As it has been shown that $\gamma=1$ and $\sigma=K$ to give the fastest guaranteed convergence Smith et al. (2018), the CoCoA is simplified to the update rule (3) that involve iterative primal and dual variables $v$ and $w$, denoted as CoCoA-PD for abbreviation.
$v^{(t+1)}_{[k]}=\underset{v_{[k]}\in\mathbb{R}^{n_{k}}}{\arg\min}\;\frac{1}{n}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(v_{i})-\frac{1}{n}\langle X_{[k]}^{\top}w^{(t)},v_{[k]}\rangle$
[…]
$w^{(t+1)}=-\frac{1}{n\lambda}\sum_{k=1}^{K}X_{[k]}v^{(t+1)}_{[k]}.$ (3)
@@ 2.2 Global Consensus ADMM with Regularization
Global consensus ADMM with regularization transforms the problem (1) into an equivalent form
$\underset{w\in\mathbb{R}^{d}}{\min}\;\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{P}_{k}}\ell_{i}(w_{k}^{\top}x_{i})+g(w)\mbox{ s.t. }w_{k}=w,k\in[K].$
and apply the standard ADMM algorithm to solve it in a distributed manner (see Section 7.1.1 of (Boyd et al., 2011)):
$\begin{split}w_{k}^{(t+1)}=&\underset{w_{k}\in\mathbb{R}^{d}}{\arg\min}\;\frac{1}{n}\sum_{i\in\mathcal{P}_{k}}\ell_{i}(w_{k}^{\top}x_{i})-\langle u_{k}^{(t)},w_{k}-w^{(t)}\rangle\\ &+\frac{\beta}{2}\|w_{k}-w^{(t)}\|^{2},k\in[K](\text{in parallel}),\\ \\ u_{k}^{(t+1)}=&u_{k}^{(t)}-\beta(w_{k}^{(t+1)}-w^{(t)}),k\in[K](\text{in parallel}),\\ w^{(t+1)}=&\underset{w\in\mathbb{R}^{d}}{\arg\min}\;g(w)-\sum_{k=1}^{K}\langle u_{k}^{(t+1)},w_{k}^{(t+1)}-w\rangle\\ &+\sum_{k=1}^{K}\frac{\beta}{2}\|w_{k}^{(t+1)}-w\|^{2}.\end{split}$
In the above algorithm, $\beta$ denotes the tuning parameter. This algorithm can be cast into an iterative update rule Consensus-PD of the primal variable $w$ and the dual variable $v$, summarized in Proposition 2.1.
Proposition 2.1.
The global consensus ADMM with regularization for solving the primal problem (1) is equivalent to the following update rule:
$v^{(t+1)}_{[k]}=\underset{v_{[k]}\in\mathbb{R}^{n_{k}}}{\arg\min}\;\frac{1}{n}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(v_{i})+\frac{1}{2n^{2}\beta}\Big{\|}v_{[k]}-v_{[k]}^{(t)}\Big{\|}_{X_{[k]}^{\top}X_{[k]}}^{2}$
[…]
$w^{(t+1)}=\mathrm{prox}_{(\beta K)^{-1}g}\left(w^{(t)}-\frac{1}{n\beta K}X\left(2v^{(t+1)}-v^{(t)}\right)\right).$ (4)
To leverage the closed-form proximal operators of $\ell_{i}^{*}(\cdot)$s in $v$-steps, the linearized ADM approach (Lin et al., 2011) can be employed, resulting in LinConsensus-PD update rule:
$v_{[k]}^{(t+1)}$ $=\mathrm{prox}_{(n\beta/\tau)\ell^{*}_{[k]}}\left(v_{[k]}^{(t)}+\frac{n\beta}{\tau}X_{[k]}^{\top}w^{(t)}\right),$
[…]
$w^{(t+1)}$ $=\mathrm{prox}_{(\beta K)^{-1}g}\left(w^{(t)}-\frac{1}{n\beta K}X\left(2v^{(t+1)}-v^{(t)}\right)\right).$ (5)

### v2

# CoCoA Is ADMM: Unifying Two Paradigms in Distributed Optimization

@@ Abstract
We consider primal-dual algorithms for general empirical risk minimization problems in distributed settings, focusing on two prominent classes of algorithms. The first class is the communication-efficient distributed dual coordinate ascent (CoCoA), derived from the coordinate ascent method for solving the dual problem. The second class is the alternating direction method of multipliers (ADMM), including consensus ADMM, proximal ADMM, and linearized ADMM. We demonstrate that both classes of algorithms can be transformed into a unified update form that involves only primal and dual variables. This discovery reveals key connections between the two classes of algorithms: CoCoA can be interpreted as a special case of proximal ADMM for solving the dual problem, while consensus ADMM is equivalent to a proximal ADMM algorithm. This discovery provides insight into how we can easily enable the ADMM variants to outperform the CoCoA variants by adjusting the augmented Lagrangian parameter. We further explore linearized versions of ADMM and analyze the effects of tuning parameters on these ADMM variants in the distributed setting. Extensive simulation studies and real-world data analysis support our theoretical findings.
@@ 1 Introduction
In this paper, we consider algorithms for solving a distributed learning problem, where $K$ machines collaboratively solve a general empirical risk minimization (ERM) problem using $n$ data samples $\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$, which are partitioned across the machines. The standard formulations of the general distributed ERM problems are given by:
$\underset{w\in\mathbb{R}^{d}}{\min}\;\mathcal{P}(w):=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{P}_{k}}\ell_{i}(w^{\top}x_{i})+g(w),$ (P)
where $\{\mathcal{P}_{k}\}_{k=1}^{K}$ denotes a partition of the dataset across the $K$ machines with $|\mathcal{P}_{k}|=n_{k}$ and $\sum_{k=1}^{K}n_{k}=n$. The parameter $w\in\mathbb{R}^{d}$ is the primal variable of interest. Each $\ell_{i}(\cdot)$ is a convex loss function associated with the $i$-th data sample, potentially involving the label information. The regularization term $g(\cdot)$ is convex and possibly non-smooth. The dual form of problem (P) is
$\underset{v\in\mathbb{R}^{n}}{\max}\;\mathcal{D}(v):=-\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{P}_{k}}\ell_{i}^{*}(v_{i})-g^{*}\left(-\frac{1}{n}Xv\right),$ (D)
where the parameter $v\in\mathbb{R}^{n}$ is the dual variable, and functions $\ell_{i}^{*}$ and $g^{*}$ in dual formulation are Fenchel conjugates of $f$ and $g$ respectively. The class of problems represented by (P) and (D) forms a foundational framework in statistical machine learning [22]. Common choices for loss function $\ell_{i}(\cdot)$ include the squared loss, least absolute deviation, quantile loss [11], Huber loss [9], and hinge loss for SVMs [23], while popular regularizers $g(\cdot)$ include the $\ell_{1}$ norm [21], $\ell_{2}$ norm, and elastic net [27].
@@
Over the past decades, efficient distributed algorithms have been designed to solve problems (P) and (D). One popular class of the algorithms to solve the problem (D) is communication-efficient distributed dual coordinate ascent (CoCoA) framework and its variants [24, 17, 20, 10, 5, 12]. This approach originates from adapting the dual coordinate descent method [19] for distributed scenarios. It create an upper bound of the dual objective which enables parallel update of the corresponding dual variables on individual machines, under the assumption that the penalty function of $g$ is strongly convex.
@@
Another class of methods that solve (P) includes the Alternating Direction Method of Multipliers (ADMM) and its variants [1]. For example, consensus ADMM method has been recognized as an effective way to solve distributed learning problems, which tailors the classical two-block ADMM method for distributed scenarios through duplicating many local variables. Recently, significant research has focused on developing ADMM variants to address the federated learning problems. An extension is generalized (or proximal) ADMM algorithms, for example [4], to achieve faster convergence rate for the ADMM algorithm. The reviews [25, 7, 6, 18] provide comprehensive discussions on the recent advancements of using ADMM variants in distributed optimization.
@@
In this study, we present an interesting discovery that CoCoA, consensus ADMM, and two distributed proximal ADMM algorithms [4, 3] can all be cast into the same kind of update rules that only involve the global update of primal variables and local updates of dual variables. Up to our knowledge, this is an original discovery, and the unified update rules result in (1) an easy calculation of the dual gap, (2) a novel understandings on the connections among CoCoA and ADMM-type algorithms in the existing literature, (3) a unified proof for the convergence of ADMM-type algorithms. A major outcome is that the update rule of CoCoA is identical to a special proximal ADMM algorithm with a specific selection of step size, whereas this proximal ADMM algorithm is equivalent to the consensus ADMM algorithm.
@@
    We reformulate consensus ADMM, linearized consensus ADMM, two proximal ADMM algorithms, and CoCoA for solving the distributed ERM problems (P) and (D) into a unified update form that involves only the primal variable $w$ and the dual variable $v$. In this unified form, the primal variable $w$ is updated centrally at the server, while the dual variable $v$ is updated in a distributed manner across local machines, where each machine $k$ maintains and updates its local dual block $v_{[k]}$. To preserve data privacy and reduce communication overhead, two known encoder functions $f_{k}(\cdot)$ and $g_{k}(\cdot)$ are used to encode information exchanged between the $k$-th machine and the central server.
@@
    Based on this unified formulation, we establish explicit connections among the three distributed algorithms (see Figure 1). Specifically, we show that for $\ell_{2}$-regularized empirical risk minimization (ERM), the dual variable updates in the CoCoA framework are equivalent to those in the first proximal ADMM algorithm, which is in turn equivalent to the consensus ADMM method. Furthermore, the linearized version of consensus ADMM aligns with the second proximal ADMM formulation, which enables the use of closed-form proximal operators for the loss function.
@@
    We thoroughly study the effects of the tuning parameters in both the proximal ADMM and consensus ADMM and use extensive real data experiments to verify our results.
The remainder of the paper is organized as follows. Section 2 reviews preliminaries on distributed primal-dual optimization. Section 3 introduces the five algorithms and cast them into a unified form of primal-dual update. In Section 4, we use the unified update form to evaluate the connections between algorithms. Section 5 further provides a unified proof of the convergence for all algorithms, leveraging the update form. Section 6 reports numerical experiments that validate our theoretical findings. Finally, Section 7 concludes the paper. All technical proofs are provided in the appendix.
Figure 1: Connections among distributed algorithms: under $\ell_{2}$-regularized ERM, CoCoA is equivalent to first Proximal ADMM with $\rho=\lambda^{-1}$, and Consensus ADMM is equivalent to first Proximal ADMM when $\beta K=\rho^{-1}$. Linearized consensus ADMM is equivalent to the linearized proximal ADMM.
2 Preliminaries
@@
Let $X=[x_{1},\ldots,x_{n}]\in\mathbb{R}^{d\times n}$ denote the full training data matrix, where each column $x_{i}\in\mathbb{R}^{d}$ is a feature vector. The corresponding dual variable is represented by a vector $v=[v_{1},\ldots,v_{n}]^{\top}\in\mathbb{R}^{n}$. In a distributed setting with $K$ machines, we denote by $v_{[k]}\in\mathbb{R}^{n_{k}}$ and $X_{[k]}\in\mathbb{R}^{d\times n_{k}}$ the local dual variable block and local data matrix stored on the $k$-th machine, respectively. The global data and dual variable can be expressed as block concatenations:
$X=[X_{[1]},\ldots,X_{[K]}],v=\Big[v_{[1]}^{\top},\ldots,v_{[K]}^{\top}\Big]^{\top}.$
We define the local Fenchel conjugate loss as $\ell_{[k]}^{*}(v_{[k]}):=\sum_{i\in\mathcal{P}_{k}}\ell_{i}^{*}(v_{i})$, where $\mathcal{P}_{k}$ is the index set of samples on machine $k$. For a symmetric positive semidefinite matrix $S$, the weighted norm is denoted by $\|x\|_{S}:=\sqrt{x^{\top}Sx}$, and $\lambda_{\max}(M)$ represents the largest eigenvalue of matrix $M$.
Proximal Operator and Moreau Identity.
For a convex function $f:\mathbb{R}^{m}\to\mathbb{R}$ and a scalar $\lambda>0$, the proximal operator is defined as:
$\text{prox}_{\lambda f}(v):=\arg\min_{x\in\mathbb{R}^{m}}\left(f(x)+\frac{1}{2\lambda}\|x-v\|^{2}\right).$
An important identity that we use throughout the paper is the Moreau decomposition:
$\text{prox}_{\lambda f}(v)+\lambda\,\text{prox}_{f^{*}/\lambda}(v/\lambda)=v,$
where $f^{*}$ is the Fenchel conjugate of $f$. This identity implies that if the proximal operator of $f^{*}$ is computationally tractable, then the proximal operator of $f$ can be efficiently computed as well.
Saddle-Point Reformulation.
The general ERM problem can be equivalently expressed as a saddle-point problem:
$\min_{w\in\mathbb{R}^{d}}\;\max_{v\in\mathbb{R}^{n}}\;\left\{L(w;v):=-\frac{1}{n}\sum_{i=1}^{n}\ell_{i}^{*}(v_{i})+\frac{1}{n}\langle w,Xv\rangle+g(w)\right\}.$ (SP)
Note that $\mathcal{D}(v):=\min_{w}L(w;v)$ and $\mathcal{P}(w):=\max_{v}L(w;v)$, we have the standard primal-dual property:
$\mathcal{D}(v)\leq L(w;v)\leq\mathcal{P}(w),\quad\text{and}\quad\mathcal{D}(v^{*})=L(w^{*};v^{*})=\mathcal{P}(w^{*}).$
This relation ensures that the saddle-point value characterizes both the optimal primal and dual solutions. We use the primal–dual certificate to monitor convergence:
$\text{Gap}=\mathcal{P}(w^{(t)})-\mathcal{D}(v^{(t)}),$
which measures the optimality gap between the primal and dual iterates $w^{(t)}$ and $v^{(t)}$ at round $t$. A smaller gap indicates that the iterates are closer to saddle-point optimality.
@@ 3 Distributed Algorithms via Primal and Dual Updates
In this section, we demonstrate that a variety of distributed algorithms—including the CoCoA algorithm with ridge regularization [10, 17, 16, 20], the global consensus ADMM algorithm [1] and its linearized variant [13], as well as two proximal ADMM methods [4]—can all be cast into a unified update framework involving only the primal and dual variables. As we will show in the following section, this unified formulation reveals important structural connections among these different techniques.
@@ 3.1 Global Consensus ADMM with Regularization
Consensus ADMM with regularization reformulates the original problem (P) into the equivalent form:
$\underset{w\in\mathbb{R}^{d}}{\min}\;\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{P}_{k}}\ell_{i}(w_{k}^{\top}x_{i})+g(w)\quad\text{s.t.}\quad w_{k}=w,\;\forall k\in[K],$
and solves it in a distributed fashion using the standard ADMM scheme (see Section 7.1.1 of [1]):
$\begin{split}w_{k}^{(t+1)}=&\arg\min_{w_{k}\in\mathbb{R}^{d}}\;\frac{1}{n}\sum_{i\in\mathcal{P}_{k}}\ell_{i}(w_{k}^{\top}x_{i})-\langle u_{k}^{(t)},w_{k}-w^{(t)}\rangle+\frac{\beta}{2}\|w_{k}-w^{(t)}\|^{2},\quad\forall k\in[K],\\ u_{k}^{(t+1)}=&\;u_{k}^{(t)}-\beta(w_{k}^{(t+1)}-w^{(t)}),\quad\forall k\in[K],\\ w^{(t+1)}=&\arg\min_{w\in\mathbb{R}^{d}}\;g(w)-\sum_{k=1}^{K}\langle u_{k}^{(t+1)},w_{k}^{(t+1)}-w\rangle+\sum_{k=1}^{K}\frac{\beta}{2}\|w_{k}^{(t+1)}-w\|^{2}.\end{split}$
Here, $\beta>0$ denotes the augmented Lagrangian parameter. This algorithm can be cast into an iterative update rule Consensus-PD of the primal variable $w$ and the dual variable $v$, summarized in Proposition 1.
Proposition 1.
The consensus ADMM with regularization for solving the primal problem (P) is equivalent to the following update rule:
$w^{(t)}$ $=\mathrm{prox}_{(\beta K)^{-1}g}\left(w^{(t-1)}-\frac{1}{n\beta K}X\left(2v^{(t)}-v^{(t-1)}\right)\right),$ (1)
$v^{(t+1)}_{[k]}$ $=\underset{v_{[k]}\in\mathbb{R}^{n_{k}}}{\arg\min}\;\frac{1}{n}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(v_{i})+\frac{1}{2n^{2}\beta}\left\|v_{[k]}-v_{[k]}^{(t)}\right\|^{2}_{X_{[k]}^{\top}X_{[k]}}-\frac{1}{n}\left\langle X_{[k]}^{\top}w^{(t)},v_{[k]}\right\rangle,\quad k\in[K].$
To simplify the updates in $v$-steps, the linearized ADMM approach [13] can be employed, resulting in LinConsensus-PD update rule:
$w^{(t)}$ $=\mathrm{prox}_{(\beta K)^{-1}g}\left(w^{(t-1)}-\frac{1}{n\beta K}X\left(2v^{(t)}-v^{(t-1)}\right)\right),$ (2)
$v_{[k]}^{(t+1)}$ $=\mathrm{prox}_{(n\beta/\tau)\ell^{*}_{[k]}}\left(v_{[k]}^{(t)}+\frac{n\beta}{\tau}X_{[k]}^{\top}w^{(t)}\right),\quad k\in[K],$